OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 24 cm?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

half the diameter is the radius what is half of 24?

OpenStudy (anonymous):

12 and hi!

OpenStudy (misty1212):

then the area is \(\pi r^2\) so \[\pi\times 12^2\]

OpenStudy (misty1212):

or if you prefer \[144\pi\]

OpenStudy (anonymous):

try starting off with \[Area = \pi*r^2\]

OpenStudy (anonymous):

where pi is 3.14

OpenStudy (anonymous):

and r is half of the diameter (the diameter is 24)

OpenStudy (anonymous):

so this is the equation so far 3.14*r^2

OpenStudy (anonymous):

452.16?

OpenStudy (anonymous):

there you go. Easy enough right?

OpenStudy (misty1212):

careful here dear!!

OpenStudy (misty1212):

"exact" value means "exact" not a decimal approximation

OpenStudy (misty1212):

don't go with the decimal, go with the "exact" answer \[144\pi\]

OpenStudy (anonymous):

but it is exact...

OpenStudy (anonymous):

try it on a calculator, that's what you get.

OpenStudy (misty1212):

no dear a decimal is only an approximation for \(\pi\) it is not exact

OpenStudy (misty1212):

calculators do not give exact answers, they give decimals

OpenStudy (anonymous):

ok, then you may take it from here. *Hands baton off to @misty1212 *

OpenStudy (misty1212):

\[3.14\] is not \(\pi\) it is an approximation of \(\pi\) to only two decimal places

OpenStudy (anonymous):

A. 576 cm2 B. 144 cm2 C. 48 cm2 D. 24 cm2 Those are the answer choices

OpenStudy (misty1212):

are there no \(\pi\)'s in those answers?

OpenStudy (anonymous):

well yes but they just didnt show up

OpenStudy (misty1212):

then go with \[144\pi\]

OpenStudy (misty1212):

those \(\pi\)'s don't show sometimes

OpenStudy (anonymous):

Yea so 144 is the right answer?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

ok I will give u a medal and can you give @okayd a medal because they helped. :) plz

OpenStudy (misty1212):

sure, just did it !

OpenStudy (anonymous):

Thanks :)